# Blockchain-empowered Federated Learning: Benefits, Challenges, and Solutions

### v1

@@ 3.1 Decentralization Architecture
Figure 5: Two decentralized architectures of the Blockchain-Empowered Federated Learning (BC-FL) system. In the completely decentralized BC-FL architecture, the client assumes the responsibility of both model training and blockchain system maintenance. However, in the partially decentralized BC-FL architecture, specialized miners are responsible for maintaining the blockchain while the client solely focuses on model training.
To address the centralization problem of the FL system, the BC-FL system connects some nodes through the blockchain to perform the same function as the FL parameter server. According to existing studies, we can examine the BC-FL system from three perspectives: consensus mechanism, degree of decentralization, and smart contracts. The consensus mechanism plays a crucial role in determining how data is recorded and verified in the blockchain and is a critical component of the BC-FL system. The PoW consensus mechanism is utilized in part of the BC-FL system, in which nodes compete to solve a mathematical puzzle. The node that solves the puzzle first is responsible for aggregating the model and training information generated in that round, which is then packed as a new block. Other blockchain nodes verify the correctness of the newly generated block. If the majority of nodes confirm that the block is correct, the new block is added to the blockchain. Section 4.1 will elaborate on additional consensus mechanisms employed in the BC-FL system, in addition to PoW.
The BC-FL system has two types of decentralization architectures based on whether all nodes are likely to participate in the blockchain consensus process: complete decentralization architecture and partial decentralization architecture. In the complete decentralized architecture (as shown in Fig. 5(a)), all participating nodes engage in both the blockchain consensus process and FL training task. This architecture places higher demands on the computing power and storage capacity of all nodes in the system. In contrast, the partially decentralized architecture (as shown in Fig. 5(b)) only allows some nodes to run the blockchain system while others only perform FL training. While less open than its completely decentralized counterpart, partially decentralized architecture offer greater efficiency gains. When constructing a BC-FL system, selecting the appropriate level of decentralization depends on the specific operational needs of the system.
3.2 Workflow of BC-FL System
In this section, we introduce Blockchain-Empowered Federated Learning (BC-FL) systems and outline the general framework of the BC-FL system. The overall structure and workflow of the BC-FL system are illustrated in Figure 6.
@@ 3.2 Workflow of BC-FL System
Figure 6: Overall structure and workflow of the blockchai-empowered federated learning (BC-FL) system.
The primary workflow of a BD-FL system can be summarized as follows:
@@
Step 1. Initialization: Each client performs environment initialization based on pre-negotiated parameters, including initializes the model structure, model parameters, and training parameters. Blockchain technology can facilitate the negotiation process among clients by storing initialization parameters on the chain. Smart contracts on the blockchain can be leveraged to coordinate and streamline the initialization process.
Step 2: Local model training. After initialization, each client trains the assigned ML model using his local dataset following the agreed-upon training method. When the local model training end condition is met (typically after reaching a certain training sequence or duration), the local training step concludes and proceeds to the next stage.
Step 3: Local model upload. In this step, each client uploads his trained local model to the blockchain system. Clients initiate transactions to upload their local models, and due to the large size of the models, clients may upload only a reference link to access the local models. A more detailed discussion of this approach will be discussed in Section 4.
@@
Step 4: Transaction broadcast. Upon receiving a transaction submission, the blockchain node disseminates it to the entire blockchain network for cross-validation. The node verifies the model according to predefined rules, and if the model is valid, it is added to the local transaction pool.
@@
Step 5: Block generation. The blockchain selects the node responsible for generating blocks in the current round based on the consensus protocol. The selected node aggregates the local models received to update the global model, which is then packaged along with the local models to generate a block.
Step 6: Global model sharing and block broadcast. The blockchain system broadcasts the newly generated blocks to share the updated global model to all clients. Verification nodes conduct a series of checks upon receiving a block to determine its validity based on specific rules. If a block successfully passes the verification process and receives confirmation from the majority of nodes, it is appended to their respective local blockchains, thus achieving network-wide consensus.
Step 7: New iteration training. Clients initiate requests to download the most recent global model from the blockchain system. Subsequently, they commence a new iteration of local model training, effectively returning to Step 2 of the process.
@@
3.3 Smart Contracts
Smart contracts enhance transparency in the BC-FL system. Operations performed by parameter servers in the traditional FL, such as model aggregation, can be deployed on the blockchain as smart contracts that run automatically. This feature directly eliminates the possibility of malicious nodes failing to adhere to rules while aggregating global models. The BC-FL system resolves the centralization problem inherent in the FL system and creates additional opportunities for FL applications in the real world. Our compilation of representative BC-FL systems can be found in Table 1.
Table 1: BC-FL Systems Based on Blockchain and Federated Learning.
Ref.
Smart contract
Degree of decentralization
Consensus algorithm
Platform
Abdel [38]
$\checkmark$
Complete
Algorand
Other
Fang [39]
$\times$
Partial
Algorand
Other
Feng [40]
$\checkmark$
Partial
PBFT/Raft
Hyperledger Fabric
Guo [41]
$\checkmark$
Partial
PBFT
Hyperledger Fabric
Jiang [42]
$\times$
Complete
DPoS
Other
Liu [43]
$\times$
Partial
PoW+PoA
Other
Lu [44]
$\times$
Complete
DPoS
Other
Nguyen [45]
$\times$
Partial
PoR
Other
Nguyen [46]
$\times$
Partial
PoW
Other
Qi [47]
$\checkmark$
Partial
Hyperledger Fabric
Qi [48]
$\checkmark$
Complete
Modified PBFT
Ethereum
Qu [49]
$\times$
Complete
PoW
Other
Rehman [50]
$\checkmark$
Complete
Ethereum
Wu [51]
$\times$
Complete
PoW
Other
Xu [52]
$\times$
Complete
Other
Xu [53]
$\checkmark$
Complete
Other
Zhang [54]
$\times$
Partial
PoW
Other
Zhao [55]
$\times$
Partial
PBFT
Other
Firstly, we organized BC-FL systems based on a completely decentralized architecture. In [53], Xu et al. proposed a BC-FL framework named Blockchain Empowered Secure and Incentive Federated Learning (BESIFL). BESIFL enables any node in the network to initiate FL training requirements. Upon receipt of a requirement, BESIFL selects computing nodes with high computation reputation scores to form a computing pool and assigns them the task of model training. Meanwhile, BESIFL chooses verification nodes with high verification reputation scores to form a verification pool and assigns them the task of model aggregation and verification using pre-defined procedures specified by the smart contract. Li et al. also proposed a completely decentralized BF-FL system, where each client acts as both a FL trainer and a blockchain miner [56]. After training their local models, clients initiate blockchain transaction requests and broadcast their models by attaching them to the transaction information. Each client aggregates the global model locally after receiving local models from all other clients and starts mining. The winning miner broadcasts a block containing global model information, which is verified by other clients and then written into the blockchain. However, this system assumes that all clients possess equal computational power, which may not be realistic in practice. In addition to the two aforementioned decentralized methods, Qu et al. designed a novel approach that utilizes a rotation mechanism with randomness to select committee members for participating in blockchain consensus [57]. This proposed blockchain consensus mechanism greatly reduces additional consumption generated by the blockchain consensus process compared to the PoW mechanism. Committee members are only responsible for aggregating and validating the global model and do not participate in training. The global model is generated by committee members and stored in the blockchain after verification. While the rotation mechanism ensures the mobility of committee members, it can ensure some level of system security. However, this consensus mechanism is only applicable in situations where the number of malicious nodes is small.
@@ 3.4 Reputation Evalutation Mechanism
    The blockchain acts as a reliable third-party ledger in the BC-FL system to document crucial information regarding each node’s reputation, including but not limited to its reputation value and various calculation bases.
@@
    In the BC-FL system, the reputation computation process can be deployed on the blockchain through a specialized reputation calculation smart contract. This approach serves to ensure both transparency and automation throughout the entire computation process, thereby guaranteeing dependable and consistent outcomes.
@@
Table 2: Blockchain-based reputation mechanism in BC-FL Systems.
Ref. Reputation Source Blockchain Usage Reputation Usage
Aggregration
Other Workers
Blockchain
Usage 1
Usage 2
Model Aggregation
Node Selection
Reward or Punishment
Chen [62]
$\checkmark$
$\checkmark$
$\checkmark$
Gao [63]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Guo [41]
$\checkmark$
$\checkmark$
Haddaji[64]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
He [58]
$\checkmark$
$\checkmark$
Kang [61]
$\checkmark$
$\checkmark$
$\checkmark$
Liu [43]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Qi [47]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Qiu [65]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Rahman [66]
$\checkmark$
$\checkmark$
$\checkmark$
Xu [52]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Zhao [67]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
@@ 3.5 Incentive Evaluation Mechanism
Ref.
Incentives
Incentive Basis
Smart Contract
Abdel [38]
Manufacturer’s discount
Reputation
$\checkmark$
Chen [72]
Token
Model accuracy
$\checkmark$
Gao [63]
Token
Model accuracy, reputation
$\checkmark$
Li [73]
Token
Model accuracy
Liu [43]
Reputation
Model accuracy
Liu [74]
Ethereum
Model accuracy
$\checkmark$
Ma [75]
Financial incentive
Model accuracy
$\checkmark$
Qu [76]
Not mentioned
Data size
Qu [77]
Not mentioned
Computing power, local data
Rehman [50]
Token
Reputation
$\checkmark$
Wang [78]
Reputation, revenue
Reputation, shaply values, and model aggregration
$\checkmark$
Weng [79]
Token
Model accuracy, block mining
Xu [52]
Token
Model accuracy, training time
$\checkmark$
Xu [80]
Not mentioned
Model accuracy, consensus Participation
Zhang [81]
Token
Training speed, computing power, and feature extractors sharing
$\checkmark$
Zhang [82]
Ethereum
Model accuracy, data size
$\checkmark$
@@ 3.6 Security Enhancement
Ref.
Transparency
Auditability
Anti-malicious nodes
Traceability
Immutability
Anti-single point of failure
Awan [83]
$\checkmark$
$\checkmark$
$\checkmark$
Cheng [84]
$\checkmark$
Feng [85]
$\checkmark$
$\checkmark$
Jia [86]
$\checkmark$
$\checkmark$
Li [73]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Liu [87]
$\checkmark$
Lu [88]
$\checkmark$
Miao [89]
$\checkmark$
$\checkmark$
$\checkmark$
$\checkmark$
Mothukuri [90]
$\checkmark$
$\checkmark$
Sun [91]
$\checkmark$
$\checkmark$
$\checkmark$
Xu [52]
$\checkmark$
$\checkmark$
$\checkmark$
Zhang [81]
$\checkmark$
$\checkmark$
Zhang [54]
$\checkmark$
$\checkmark$
$\checkmark$
@@
Anti-single point of failure: The term ”single point of failure” refers to a scenario where a sole parameter server becomes the bottleneck for FL security, rendering the entire system inoperable if it fails due to an attack or power outage, among other reasons. To tackle this problem, the BC-FL system replaces the role of the parameter server with blockchain technology. As discussed in Chapter 3.1, the issue of single point of a failure is elaborated upon.
@@
In [73], Li et al. proposed a Byzantine-resistant consensus mechanism named Proof of Accuracy, which serves to identify models of poor quality. This consensus algorithm takes into consideration not only the exclusion of local models that are deemed too poor for aggregation into the global model but also the potential for a local model with a high loss value to aid the global model in escaping local optimal solutions. To fulfil this requirement, the consensus algorithm employs two critical thresholds: the accuracy oscillation threshold (AOT) and the accuracy deviation threshold (ADT). The AOT determines the maximum acceptable accuracy reduction permitted by the accepted model, while the ADT determines the maximum absolute difference in accuracy among different client models. These two thresholds are subject to dynamic adjustments as the algorithm progresses. In [65], Qiu et al. increased the security of the BC-FL system through the introduction of a novel consensus protocol called Proof of Learning (PoL). In contrast to PoW, PoL requires nodes to compete for the privilege of accounting rights through calculation by training a FL model, where the node with the smallest loss value adds a new block as the winner. Other clients aggregate the winner’s local model based on the reputation value against the winning node after verifying the authenticity of the newly added block.
@@ 4.1 Efficiency Challenges and Solutions
The decentralized nature inherent in blockchain systems presents challenges to their efficiency. For instance, consider Bitcoin, which is limited to processing only 7 transactions per second [92]. Additionally, it often necessitates waiting for several consecutive blocks before a transaction can be deemed immutable. Consequently, this leads to confirmation times extending to tens of minutes for transactions. In contrast, centralized payment systems have the capability to handle thousands of transactions per second, thereby facilitating real-time payments [93]. The primary efficiency challenges along with their corresponding solutions within the BC-FL systems are visualized in Fig. 9.
Figure 9: Efficiency challenges and related solutions in the BC-FL systems. The efficiency of blockchain is susceptible to factors such as network and computing overhead. Consequently, BC-FL systems potentially lead to a decrease in overall efficiency. In response to thus challenges, multiple strategies are contributed to mitigate the reduction in system efficiency. These methods include but not limited to, the implementation of more streamlined consensus mechanisms, reinforcement learning methods, and the refinement of blockchain topologies.
As shown in Fig. 9, while adjusting the block generation time can significantly bolster the consensus efficiency of the BC-FL system when compared to Bitcoin, blockchain involves supplementary processes like verification, communication, and network-wide consensus to ensure its functionality. Consequently, the BC-FL system unavoidably results in a reduction of overall system efficiency. To counteract the efficiency challenges stemming from blockchain integration, existing research employs a range of strategies, such as efficient consensus mechanisms, reinforcement learning, and optimized blockchain topologies. A comprehensive comparison between different solutions for each efficiency challenge is presented in Table 5.
@@
Ref.
Solutions
Used methods
Dataset
Evaluation indicators
Cao [94]
Blockchain topology
DAG blockchain
MNIST
Accuracy, loss, iteration delay
Cheng [84]
Consensus algorithm, blockchain
Two-layer blockchain, Raft, PBFT
Latency reduction
Feng [85]
Consensus algorithm, blockchain
Two-layer blockchain, sharding
MNIST
Accuracy, time cost
Hieu [95]
RL
DRL
Energy consumption, latency, total payment
Li [96]
Consensus algorithm
committee consensus mechanism
FEMNIST
accuracy, communication overhead
Lu [44]
RL, blockchain
DAG blockchain, DRL
Accuracy, time cost, agent reward, cumulative cost
Nguyen [45]
Consensus algorithm
Proof of reputation
DarkCOVID, ChestCOVID
Running latency, Block verification latency, Accuracy, Loss
Nguyen [46]
RL
DRL+A2C
SVHN, Fashion-MNIST
Accuracy, agent reward, latency, Loss
Qi [48]
Consensus algorithm
Modified PBFT
Diabetes Breast Cancer
Accuracy, time cost, gas cost
Qu [77]
Consensus algorithm
Proof of federalism
CIFAR-10
Accuracy
Xu [80]
Consensus algorithm, blockchain
Two-layer blockchain, proof of credit, efficient BFT
MNIST
Latency, communication overhead, data throughput,
Zhao [55]
RL
Federated DDQL
Agent reward, latency
4.1.1 Consensus Algorithms

### v2

@@ 3.1 Decentralization
Figure 5: Two decentralized architectures of the Blockchain-Empowered Federated Learning system.
Next, we will examine some representative architectures of completely decentralized BC-FL systems.
In [63], Xu et al. proposed a BC-FL framework named Blockchain Empowered Secure and Incentive Federated Learning (BESIFL). BESIFL enables any node in the network to initiate FL training requirements. Upon receipt of a requirement, BESIFL selects computing nodes with high computation reputation scores to form a computing pool and assigns them the task of model training. Meanwhile, BESIFL chooses verification nodes with high verification reputation scores to form a verification pool and assigns them the task of model aggregation and verification using pre-defined procedures specified by the smart contract. Li et al. also proposed a completely decentralized BF-FL system, where each client acts as both a FL trainer and a blockchain miner [72]. After training their local models, clients initiate blockchain transaction requests and broadcast their models by attaching them to the transaction information. Each client aggregates the global model locally after receiving local models from all other clients and starts mining. The winning miner broadcasts a block containing global model information, which is verified by other clients and then written into the blockchain. However, this system assumes that all clients possess equal computational power, which may not be realistic in practice. In addition to the two aforementioned decentralized methods, Qu et al. designed a novel approach that utilizes a rotation mechanism with randomness to select committee members for participating in blockchain consensus [73]. This proposed blockchain consensus mechanism greatly reduces additional consumption generated by the blockchain consensus process compared to the PoW mechanism. Committee members are only responsible for aggregating and validating the global model and do not participate in training. The global model is generated by committee members and stored in the blockchain after verification. While the rotation mechanism ensures the mobility of committee members, it can ensure some level of system security. However, this consensus mechanism is only applicable in situations where the number of malicious nodes is small.
Workflow of BC-FL Systems. The overall workflow of the BC-FL system is illustrated in Fig. 6. Different BC-FL systems may be adjusted according to specific circumstances. The steps are explained as follows:
Figure 6: Overall structure and workflow of the blockchain-empowered federated learning system.
@@
Step 1. Initialization: Each client initializes the environment based on prior negotiation, including model parameter, and training parameter. The blockchain can assist clients in negotiation by storing initialization parameters on the chain and using smart contracts.
Step 2: Local model training. Each client trains the global model using their local dataset.
Step 3: Local model upload. Clients upload training-related data and local models to the blockchain system. To alleviate storage pressure on the blockchain, clients may upload only model-related information rather than the entire model, as detailed in Section 4.3.
@@
Step 4: Transaction broadcast. Upon receiving the transaction, blockchain nodes broadcast it within the system for cross-validation. The nodes inspect the transaction content (e.g., model) based on pre-defined rules. If no issues are found, the transaction is added to the local transaction pool.
@@
Step 5: Block generation. Blockchain nodes select the node with the right to generate blocks for the current round based on the consensus protocol. This node aggregates the local models to generate the global model, compiles relevant model and training information, and creates a block.
Step 6: Block Broadcast. The blockchain system broadcasts the newly generated block. Upon receiving it, validation nodes verify the block according to specific rules. If the majority of nodes validate the block, it is added to their locally maintained blockchain, achieving consensus across the network.
Step 7: Global Model Download. Clients download the latest global model from the blockchain system.
@@ 3.2 Reputation Evalutation Mechanism
    The blockchain acts as a reliable third-party ledger in the BC-FL system to document crucial information regarding each node’s reputation, including but not limited to its reputation value [51, 75, 76] and various calculation bases [77, 53, 57].
@@
    In the BC-FL system, the reputation computation process can be deployed on the blockchain through a specialized reputation calculation smart contract [77, 78, 57]. This approach serves to ensure both transparency and automation throughout the entire computation process, thereby guaranteeing dependable and consistent outcomes.
@@
Table 2: Blockchain-based reputation mechanism in BC-FL Systems.
Ref. Reputation Source Blockchain Usage Reputation Usage
Aggregration Other Workers Blockchain Usage 1 Usage 2 Model Aggregation Node Selection Reward or Punishment
Chen [80] $\checkmark$ $\checkmark$ - - - - $\checkmark$ -
Gao [81] $\checkmark$ - - $\checkmark$ $\checkmark$ - $\checkmark$ $\checkmark$
Guo [51] $\checkmark$ - - - - - $\checkmark$ -
Haddaji[76] $\checkmark$ - $\checkmark$ $\checkmark$ - - $\checkmark$ -
He [74] $\checkmark$ - - - - - $\checkmark$ -
Kang [77] $\checkmark$ - - $\checkmark$ - - $\checkmark$ -
Liu [53] $\checkmark$ - - $\checkmark$ - - $\checkmark$ $\checkmark$
Qi [57] $\checkmark$ $\checkmark$ - $\checkmark$ $\checkmark$ $\checkmark$ - $\checkmark$
Qiu [78] $\checkmark$ - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ - -
Rahman [82] $\checkmark$ - - $\checkmark$ - - $\checkmark$ -
Xu [62] $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhao [83] $\checkmark$ - - $\checkmark$ - - $\checkmark$ $\checkmark$
Wang [66] $\checkmark$ - $\checkmark$ $\checkmark$ - $\checkmark$ - -
Lin [84] $\checkmark$ - $\checkmark$ $\checkmark$ - - $\checkmark$ -
Fu [85] $\checkmark$ - - $\checkmark$ - - - $\checkmark$
Wahrstatter [71] $\checkmark$ - - $\checkmark$ $\checkmark$ - - $\checkmark$
@@ 3.3 Incentive Evaluation Mechanism
Ref. Incentives Incentive Basis Smart Contract
Abdel [48] Manufacturer’s discount Reputation $\checkmark$
Chen [90] Token Model accuracy $\checkmark$
Gao [81] Token Model accuracy, reputation $\checkmark$
Li [91] Token Model accuracy -
Liu [53] Reputation Model accuracy -
Liu [11] Ethereum Model accuracy $\checkmark$
Ma [92] Financial incentive Model accuracy $\checkmark$
Qu [93] Not mentioned Data size -
Qu [94] Not mentioned Computing power, local data -
Rehman [60] Token Reputation $\checkmark$
Wang [95] Reputation, revenue Reputation, shaply values, and model aggregration $\checkmark$
Weng [96] Token Model accuracy, block mining -
Xu [62] Token Model accuracy, training time $\checkmark$
Xu [97] Not mentioned Model accuracy, consensus Participation -
Zhang [98] Token Training speed, computing power, and feature extractors sharing $\checkmark$
Zhang [99] Ethereum Model accuracy, data size $\checkmark$
Wang [66] Token Model accuracy, block mining -
He [100] Token Model accuracy, node behavior $\checkmark$
Wahrstatter [71] Ethereum Model accuracy, reputation $\checkmark$
@@ 3.4 Security Enhancement
Anti-single point of failure: The term "single point of failure" refers to a scenario where a sole parameter server becomes the bottleneck for FL security, rendering the entire system inoperable if it fails due to an attack or power outage, among other reasons. To tackle this problem, the BC-FL system replaces the role of the parameter server with blockchain technology. As discussed in Section 3, the issue of single point of a failure is elaborated upon.
Ref. Transparency Auditability Anti-malicious nodes Traceability Immutability Anti-single point of failure
Awan [101] - $\checkmark$ $\checkmark$ - $\checkmark$ -
Cheng [102] - - $\checkmark$ - - -
Feng [103] - - $\checkmark$ - $\checkmark$ -
Jia [104] - $\checkmark$ - - $\checkmark$ -
Li [91] - $\checkmark$ $\checkmark$ $\checkmark$ - $\checkmark$
Liu [105] - - $\checkmark$ - - -
Lu [106] - - - - - $\checkmark$
Miao [107] $\checkmark$ - $\checkmark$ $\checkmark$ - $\checkmark$
Mothukuri [108] - $\checkmark$ - - $\checkmark$ -
Sun [109] - $\checkmark$ $\checkmark$ - $\checkmark$ -
Xu [62] $\checkmark$ - - - $\checkmark$ $\checkmark$
Zhang [98] $\checkmark$ - - - $\checkmark$ -
Zhang [64] - - - $\checkmark$ $\checkmark$ $\checkmark$
Huang [110] - - - - $\checkmark$ -
Ouyang [67] - $\checkmark$ $\checkmark$ - $\checkmark$ $\checkmark$
Lin [84] - - $\checkmark$ - - $\checkmark$
Aloqaily [69] - - $\checkmark$ - - -
Mu [70] - - $\checkmark$ - - $\checkmark$
Fu [85] $\checkmark$ - - - $\checkmark$ $\checkmark$
He [100] - $\checkmark$ $\checkmark$ - $\checkmark$ $\checkmark$
Xu [111] - - $\checkmark$ $\checkmark$ $\checkmark$ -
@@
In [91], Li et al. proposed a Byzantine-resistant consensus mechanism named Proof of Accuracy, which serves to identify models of poor quality. This consensus algorithm takes into consideration not only the exclusion of local models that are deemed too poor for aggregation into the global model but also the potential for a local model with a high loss value to aid the global model in escaping local optimal solutions. To fulfil this requirement, the consensus algorithm employs two critical thresholds: the accuracy oscillation threshold (AOT) and the accuracy deviation threshold (ADT). The AOT determines the maximum acceptable accuracy reduction permitted by the accepted model, while the ADT determines the maximum absolute difference in accuracy among different client models. These two thresholds are subject to dynamic adjustments as the algorithm progresses. In [78], Qiu et al. increased the security of the BC-FL system through the introduction of a novel consensus protocol called Proof of Learning (PoL). In contrast to PoW, PoL requires nodes to compete for the privilege of accounting rights through calculation by training a FL model, where the node with the smallest loss value adds a new block as the winner. Other clients aggregate the winner’s local model based on the reputation value against the winning node after verifying the authenticity of the newly added block. Ouyang et al. utilized smart contracts to authenticate participating nodes and prevent malicious nodes from participating [67].
@@ 4.1 Efficiency Challenges and Solutions
The processing capacity of blockchain systems is inherently limited. For instance, Bitcoin can only handle seven transactions per second [112]. In contrast, modern centralized payment systems can process thousands of transactions per second [113]. Fig. 9 illustrates the efficiency challenges faced by blockchain in BC-FL systems. Unlike centralized systems, blockchain systems necessitate additional steps such as verification, communication, and network-wide consensus to maintain normal operations, which reduces the efficiency when integrating blockchain with FL systems. Current BC-FL systems address these efficiency issues through various methods, including efficient consensus mechanisms, reinforcement learning (RL), and optimized blockchain topologies. A summary of the pertinent literature is provided in Table 5.
Figure 9: Efficiency challenges and related solutions in the BC-FL systems. The efficiency of blockchain is susceptible to factors such as network and computing overhead. Consequently, BC-FL systems potentially lead to a decrease in overall efficiency. In response to thus challenges, multiple strategies are contributed to mitigate the reduction in system efficiency. These methods include but not limited to, efficient consensus mechanisms, reinforcement learning, and optimized blockchain topologies.
@@
Ref. Solutions Detailed methods Dataset Evaluation indicators
Cao [114] Blockchain topology DAG blockchain MNIST Accuracy, loss, iteration delay
Cheng [102] Consensus algorithm, blockchain topology Two-layer blockchain, Raft, PBFT - Latency reduction
Feng [103] Consensus algorithm, blockchain topology Two-layer blockchain, sharding MNIST Accuracy, time cost
Hieu [115] RL DRL - Energy consumption, latency, total payment
Li [116] Consensus algorithm committee consensus mechanism FEMNIST accuracy, communication overhead
Lu [54] RL, blockchain DAG blockchain, DRL - Accuracy, time cost, agent reward, cumulative cost
Nguyen [55] Consensus algorithm Proof of reputation DarkCOVID, ChestCOVID Running latency, block verification latency, Accuracy, Loss
Nguyen [56] RL DRL, A2C SVHN, Fashion-MNIST Accuracy, agent reward, latency, Loss
Qi [58] Consensus algorithm Modified PBFT Diabetes Breast Cancer Accuracy, time cost, gas cost
Qu [94] Consensus algorithm Proof of federalism CIFAR-10 Accuracy
Xu [97] Consensus algorithm, blockchain topology Two-layer blockchain, proof of credit, efficient BFT MNIST Latency, communication overhead, data throughput
Zhao [65] RL Federated DDQL - Agent reward, latency
Wang [110] Blockchain topology Two-layer blockchain TSP, FMNIST Accuracy, energy consumption, learning utility
Yuan [68] Consensus algorithm, blockchain topology Blockchain sharding, DAG-based mainchain MNIST, Penn Treebank Accuracy, training Latency, testing perplexity
Lin [84] Blockchain topology, RL Blockchain sharding, DRL-based sharding MNIST, KMNIST, FMNIST, CIFAR-10 Accuracy, agent reward, reputation of nodes
4.1.1 Efficient Consensus Algorithms